OpenStudy (anonymous):

Suppose that x* means 5x^2 - x, then what does (y+3)* mean?

Directrix (directrix):

(y+3)* = 5y^2+30y+45 - (y + 3) (y+3)* = 5y^2+30y+45 -1 (y + 3) (y+3)* = 5y^2+30y+45 -y -3 (y+3)* = 5y^2 +29y - 42

OpenStudy (asnaseer):

The way to solve this is to replace every occurrence of 'x' in the original equation with 'y+3'.\[\begin{align} x*&\implies5x^2-x\\ \text{therefore, }(y+3)*&\implies5(y+3)^2-(y+3)\\ &\implies5(y^2+6y+9)-y-3\\ &\implies5y^2+30y+45-y-3\\ &\implies5y^2+29y+42 \end{align}\]this is how @Directrix got to the same answer.